OpenStudy (anonymous):

find an equation of the line containing the given pair of points. (-5,0) (0,9) y=

OpenStudy (acxbox22):

sure

OpenStudy (anonymous):

ok thanks

OpenStudy (acxbox22):

y=9/5x+9

OpenStudy (anonymous):

write an equation of the line containing the given point. (0,3) 2x+7y=3 @acxbox22

OpenStudy (acxbox22):

y=-2/7x+3